OpenStudy (anonymous):

a shop sells candies for 60 cents and 10 cents each junior spent $7.40 and bought 24 candies how many 60 cent candies did he buy

OpenStudy (anonymous):

x + y = 24 .6x + .1y = 7.40

OpenStudy (anonymous):

x = 24-y .6 (24-y) + .1y = 7.40 14.4 - .6y + .1y = 7.40 14.4-7.40 = .5y 7.00 = .5y y = 14 x= 24-14 = 10 he bought 10 of the .60 candies